OpenStudy (anonymous):

Which shows three ordered pairs that are all solutions of the equation y = x + 4? A. (1, 4), (2, 6), and (3, 7) B. (1, 4), (2, 8), and (3, 12) C. (3, 7), (5, 9), and (7, 10) D. (3, 7), (5, 9), and (7, 11)

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

take each option in turn and substitute the values of x and y into the equation to see which ones fit. for example A (1,4):- x = 1 and y = 4:- 4 = 1 + 4 4 = 5 its not this one so you go on to B and try the ordered pairs in B as with A if the first 1 does n't fit theres no need to try the others

OpenStudy (welshfella):

in fact , (1,4) is in B so you can do on to C

OpenStudy (anonymous):

@welshfella How does 4 = 5?.

OpenStudy (welshfella):

it doesn't of course - so that means (1,4) cannot be a solution to the equation y = x + 4

OpenStudy (welshfella):

it (1,4) was a solution we'd have the same number like 5=5

OpenStudy (anonymous):

Oh okay. so a and b is done o time for c and d.?.

OpenStudy (welshfella):

- I mean the same number would be either side of the '='

OpenStudy (welshfella):

right its either c or d

OpenStudy (welshfella):

try (3 ,7) x =3 y=7 7 = 3 + 4 7 = 7 so (3,7) is a solution

OpenStudy (welshfella):

now you need to try (5,9)

OpenStudy (anonymous):

9 = 4 + 5 9=9

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

- yes that means that (5,9) satisfies the equation (5,9) is a solution

OpenStudy (welshfella):

now see if (7,10 0 is a solution

OpenStudy (welshfella):

(7,10) - that is y=10 and x=7

OpenStudy (anonymous):

(5,9) 5 X=5,Y=9 9 = 4 + 5 9 = 9 So its a solution?.

OpenStudy (anonymous):

(7,10) 10 = 3 + 7 10 = 10 So it is a sulution @welshfella @pooja195 @Godlovesme @BlossomCake

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

So its C?..@welshfella

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

does (7,10) fit the equation ? try it

OpenStudy (welshfella):

you have to try all 3 ordered pairs

OpenStudy (anonymous):

10 = 3 + 7

OpenStudy (welshfella):

no the equation is y = x + 4 so you need to put y = 10 and x = 7 into this equation

OpenStudy (welshfella):